A rectangular piece of metal is 10 in longer than it is wide. Squares with sides
2 in long are cut from the four corners and the flaps are folded upward to form an open box. If the volume of the box is 528 in3​, what were the original dimensions of the piece of​ metal?

Answer by ankor@dixie-net.com(22740)

A rectangular piece of metal is 10 in longer than it is wide.
Squares with sides 2 in long are cut from the four corners and the flaps are folded upward to form an open box.
If the volume of the box is 528 in3​, what were the original dimensions of the piece of​ metal?
:
let x = the width of the rectangular piece of metal
then
(x+10) = the length
:
Removing the 2" squares would make the dimensions of the box:
(x-4) by (x+10-4) or (x-4) by (x+6)
the height of the box = 2"
:
The volume equation L * W * h
(x+6)*(x-4)*2 = 528
divide both sides by 2
(x+6)(x-4) = 264
FOIL
x^2 - 4x + 6x - 24 = 264
x^2 + 2x - 24 - 264 = 0
a quadratic equation
x^2 + 2x - 288 = 0
You can use the quadratic formula a=1; b=2; c=-288, but this will factor to
(x+18)(x-16) = 0
the positive solution
x = 16 in is the width of the rectangular metal
then
26 in = the length
:
:
Check this, the dimensions of the box will be 4' less and the length and width
22 * 12 * 2 = 528 cu in